Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 201     Entry time: 11 May 2023, 23:48
Author: Romain Letrun 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 9.3 keV, 1.4-1.8 mJ and 12 keV, 1.1 mJ. 202 pulses most of the day.
  • Reduced slope after afternoon tuning, down to 20%.
  • Change to 12 keV with file prepared earlier this week did not work (no lasing). Tuning to 12 keV took 40 minutes.
  • 30 minutes downtime in the morning.

Optical laser delivery

  • NA

Achievements/Observations

  • Optimised slits to protect the nozzle from X-ray damage.
  • Collected data with a dozen of user samples. High hit rate and resolution achieved with some of them.

Issues

  • No major issues.
ELOG V3.1.4-7c3fd00